OpenStudy (anonymous):

show all work. A simple dartboard has three areas… the main board has a radius of 8 inches, there is a circle with a radius of 5 inches, and the bullseye has a radius of 2 inches. What is the probability of a random dart landing inside the bullseye? Round to the nearest thousandth

OpenStudy (amistre64):

Area at r=2 ---------- i assume Area at r=8

OpenStudy (amistre64):

4/64 = 1/16

OpenStudy (amistre64):

2.5 4.25 8.125 16.0625 .0625

OpenStudy (anonymous):

thank you,but can you and will you please explain how you got the answers from 2.5 to .0625. Iam confused!

OpenStudy (amistre64):

oh... i was tryng to remember how to but fractions to decimals; i was a carpenter for many years and i would use that to convert between the two

OpenStudy (amistre64):

1/2 = .50 1/4 = .250 1/8 = .1250 1/16 = .0625

OpenStudy (anonymous):

ok thank you, i never had this math it's been 33 years out of school

OpenStudy (amistre64):

its was abt 20 yrs b4 i went to college